OpenStudy (idkwut):

Water is pouring into a container at a rate given by ri = t^3(gallons/hour), and water is pouring out of the same container at a rate given by ro(t) = e^t(gallons/hour). If there are 100 gallons of water at time t = 0, how many gallons of water will be in the container at t=4 hours?

OpenStudy (idkwut):

@ganeshie8

ganeshie8 (ganeshie8):

rate of change of water in container = (in rate) - (out rate)

ganeshie8 (ganeshie8):

say "y" is the amount of water present in container at any given time

ganeshie8 (ganeshie8):

then you can setup a differential equation : \(\large \dfrac{dy}{dt} = t^3 - e^t\)

ganeshie8 (ganeshie8):

solve it

OpenStudy (idkwut):

@ganeshie8 64 - e^4

ganeshie8 (ganeshie8):

doesnt look right

ganeshie8 (ganeshie8):

whats ur solution to above equation ?

ganeshie8 (ganeshie8):

Ahh I see where the mistake was !

OpenStudy (idkwut):

9.6, but I think that's wrong

ganeshie8 (ganeshie8):

you need to solve the differential equation first, before plugging in t = 4

ganeshie8 (ganeshie8):

\(\large \dfrac{dy}{dt} = t^3 - e^t\) Solving this means, solving the function "\(y\)"

ganeshie8 (ganeshie8):

Integrate both sides

ganeshie8 (ganeshie8):

\(\large \dfrac{dy}{dt} = t^3 - e^t\) \(\large y = \int t^3 - e^t dt\)

ganeshie8 (ganeshie8):

what do u get ?

OpenStudy (idkwut):

t^4/4 -e^t, then we plug in 4?

OpenStudy (idkwut):

I still got 9.4 lol

ganeshie8 (ganeshie8):

Yes, but not yet... where is the arbitrary constant ?

ganeshie8 (ganeshie8):

\(\large \dfrac{dy}{dt} = t^3 - e^t\) \(\large y = \int t^3 - e^t dt\) \(\large y = \dfrac{t^4}{4} - e^t + c\)

ganeshie8 (ganeshie8):

right ?

ganeshie8 (ganeshie8):

we're given : ` If there are 100 gallons of water at time t = 0,`

OpenStudy (idkwut):

Oh, yes. Lol, I always forgot the constant and my teacher marks points off for it -_-

ganeshie8 (ganeshie8):

hahah happens :) plugin t = 0, y = 100 and solve \(c\)

OpenStudy (idkwut):

100 = (0^4)/4 + (e^0) + C = 101

OpenStudy (idkwut):

Then we do it for t - 4 and subtract?

OpenStudy (idkwut):

I meant t = 4

ganeshie8 (ganeshie8):

c = 101 is correct, so the equation for amount water in tank is : \(\large y = \dfrac{t^4}{4} - e^t + 101\)

ganeshie8 (ganeshie8):

yes... simply plugin t = 4 that gives amount of water in tank after 4 hours

OpenStudy (idkwut):

110.402

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)